# On stabilizers in finite permutation groups

### v3

@@ 3.2. Groups of product type
Let $X$ be a primitive almost simple group with solvable point stabilizer $Y<X$, and let $T\lhd X$ be its socle. Let $d\geq 2$, and let $S\leqslant\mathop{\mathrm{Sym}}(d)$ be a transitive solvable group. By [LZ11, Theorem 1.1], the primitive group $G$ that we care about satisfies
@@
Let $V\cong X^{d}$ be the base subgroup of $X\wr_{d}S$, and let $H<G$ be the stabilizer of a point in the product action. Let $W\cong Y^{d}$ be the subgroup of $V$ such that $H=WS\cap G$. Stabilizing two points corresponds to intersecting $H$ with one of its conjugates. To prove the theorem, we will find $v\in G\cap V$ such that the derived length of
@@
is bounded by an absolute constant (actually $13$). The proof will follow, because $H\cap H^{v}\leqslant I_{v}$, and the derived length of a subgroup is bounded by the derived length of the group. Let $\pi\colon G\to S$ be the natural projection, and let $v\in V$. Since $Ker(\pi)=G\cap V$, we have
@@
Now $dl(Y)\leq 10$ by Theorem 1.4, and the rest of the proof deals with $\pi(I_{v})$. It is not hard to see that $C_{S}(v)=S\cap S^{v}\subseteq\pi(I_{v})$, but in general $\pi(I_{v})$ can be quite bigger. For every $w\in W$, $\sigma\in S$ and $v\in V$, an explicit computation gives
@@
This suggests to look at the double cosets of $Y$ in $X$ as colors, and connects Theorem 1.5 with the matter of Section 2. Let $c_{1},\ldots,c_{r}\in X$ be representatives of these double cosets ($r\geq 2$). For any $v\in V$, and $j\in\{1,\ldots,r\}$, let
$\Delta_{j}=\{i\in\{1,\ldots,d\}:x_{i}\in Yc_{j}Y\}.$
@@
(3.3) $\pi(I_{v})\>=\>\cap_{j=1}^{r}Stab_{S}(\Delta_{j})\>=\>\cap_{j=1}^{r-1}Stab_{S}(\Delta_{j}).$
To construct $v$, we will use only two elements $c_{1}$ and $c_{2}$ (and so two sets $\Delta_{1}$ and $\Delta_{2}=\{1,\ldots,d\}\setminus\Delta_{1}$). One reason is that we want $v$ to lie in $G\cap V$, and not just in $V$. This is assured if we choose representatives $c_{i}$’s which are in $T$. We can surely choose $c_{1}=1$ and $c_{2}\in T\setminus\{1\}$, because otherwise $T\leqslant Y$, which is impossible because $Y$ is core-free. Finally, if $\Delta_{1}=\Delta\subseteq\{1,\ldots,d\}$ is as provided by Corollary 1.2, from (3.3) we get $dl(\pi(I_{v}))\leq 3$, concluding by (3.1) that $dl(I_{v})\leq 13$ as desired. ∎
@@
In the proof of Theorem 1.5, it might seem wasteful to construct $v$ with only two elements $c_{1}$ and $c_{2}$. As explained, one reason is that we want $v\in G$, but there is also a different motivation. In fact, it is well known that the number $r$ of double cosets of $Y$ in $X$ equals the rank of the primitive group $X$ with solvable stabilizer $Y$, and there actually exist some doubly transitive groups of this type (as $\mathop{\mathrm{Sym}}(5)$, for example). A more detailed analysis of almost simple groups, and the additional use of Corollary 2.14, might help to find the right constant in Theorem 1.5. In the case of primitive solvable groups, this lies somewhere between $4$ and $9$ (see [LS24, Remark 2.9]). H.Y. Huang pointed out that $5$ is attained by the almost simple group $\mathop{\mathrm{Aut}}(\mathrm{P}\Omega^{+}_{8}(3))$. With some more work, it is not hard to use this information to construct a group of product type where the constant $6$ is achieved.
@@
Let $\delta\in\{2,3,4\}$. Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a primitive group with solvable stabilizer. Then there exists $\Delta\subseteq\Omega$, $|\Delta|=\delta$, such that $Stab_{G}(\Delta)$ is solvable of derived length bounded by an absolute constant.
@@
Let $\{x_{1},x_{2}\}\subseteq\Omega$ be as provided by Theorem 1.5. If $\{x_{1},x_{2}\}\subseteq\Delta$, then $C_{G}(\Delta)$ is solvable of bounded derived length. Now $Stab_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(4)$, and it follows that $Stab_{G}(\Delta)$ is solvable of bounded derived lenght. ∎

### v4

@@ 3.2. Groups of product type
Let $X$ be a primitive almost simple group with solvable point stabilizer $Y<X$, and let $T\lhd X$ be its socle. Let $d\geq 2$, and let $S\leqslant\mathop{\mathrm{Sym}}(d)$ be a transitive solvable group. By [LZ11, Th. 1.1], the primitive group $G$ that we care about satisfies
@@
By the discussion above, it rests only to deal with the product type case. Let $V\cong X^{d}$ be the base subgroup of $X\wr_{d}S$, and let $H<G$ be the stabilizer of a point in the product action. Let $W\cong Y^{d}$ be the subgroup of $V$ such that $H=WS\cap G$. Stabilizing two points corresponds to intersecting $H$ with one of its conjugates. To prove the theorem, we will find $v\in G\cap V$ such that the derived length of
@@
is bounded by an absolute constant (actually $13$). The proof will follow, because $H\cap H^{v}\leqslant I_{v}$, and the derived length of a subgroup is bounded by the derived length of the group. Let $\pi\colon G\to S$ be the natural projection, and let $v\in V$. Since $\mathop{\mathrm{Ker}}(\pi)=G\cap V$, we have
@@
Now $dl(Y)\leq 10$ by Theorem 1.4, and the rest of the proof deals with $\pi(I_{v})$. It is not hard to see that $C_{S}(v)=S\cap S^{v}\subseteq\pi(I_{v})$, but in general $\pi(I_{v})$ can be much bigger. For every $w\in W$, $\sigma\in S$ and $v\in V$, an explicit computation gives
@@
This suggests to look at the double cosets of $Y$ in $X$ as colors, and connects Theorem 1.5 with the matter of Section 2. Let $c_{1},\ldots,c_{r}\in X$ be representatives of these double cosets ($r\geq 2$). For any $v=(x_{1},\ldots,x_{d})\in V$, and $j\in\{1,\ldots,r\}$, let
$\Delta_{j}=\Delta_{j}(v)=\{i\in\{1,\ldots,d\}:x_{i}\in Yc_{j}Y\}.$
@@
(3.3) $\pi(I_{v})\>=\>\cap_{j=1}^{r}\mathop{\mathrm{Stab}_{S}}(\Delta_{j})\>=\>\cap_{j=1}^{r-1}\mathop{\mathrm{Stab}_{S}}(\Delta_{j}).$
To construct $v$, we will use only two elements $c_{1}$ and $c_{2}$ (and so two sets $\Delta_{1}$ and $\Delta_{2}=\{1,\ldots,d\}\setminus\Delta_{1}$). One reason is that we want $v$ to lie in $G\cap V$, and not just in $V$. This is assured if we choose representatives $c_{i}$’s which are in $T$. We can surely choose $c_{1}=1$ and $c_{2}\in T\setminus Y$, because otherwise $T\leqslant Y$, which is impossible because $Y$ is core-free. Finally, if $\Delta_{1}=\Delta\subseteq\{1,\ldots,d\}$ is as provided by Corollary 1.2, from (3.3) we get $dl(\pi(I_{v}))\leq 3$, concluding by (3.1) that $dl(I_{v})\leq 13$ as desired. ∎
@@
In the proof of Theorem 1.5, it might seem wasteful to construct $v$ with only two elements $c_{1}$ and $c_{2}$. As explained, one reason is that we want $v\in G$, but there is also a different motivation. In fact, it is well known that the number $r$ of double cosets of $Y$ in $X$ equals the rank of the primitive group $X$ with solvable stabilizer $Y$, and there actually exist various doubly transitive groups of this type. A more detailed analysis of almost simple groups, and the additional use of Corollary 2.14, might help to find the right constant in Theorem 1.5. In the case of primitive solvable groups, this lies somewhere between $4$ and $9$ [LS25, Rem. 2.9]. H.Y. Huang pointed out that $5$ is attained by the almost simple group $\mathop{\mathrm{Aut}}(\mathrm{P}\Omega^{+}_{8}(3))$. With some more work, it is not hard to use this information to construct a group of product type where the constant $6$ is achieved.
@@
Let $\delta\in\{2,3,4\}$. Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a primitive group with solvable stabilizer. Then there exists $\Delta\subseteq\Omega$, $|\Delta|=\delta$, such that $\mathop{\mathrm{Stab}}_{G}(\Delta)$ is solvable of derived length bounded by an absolute constant.
@@
Let $\{x_{1},x_{2}\}\subseteq\Omega$ be as provided by Theorem 1.5. If $\{x_{1},x_{2}\}\subseteq\Delta$, then $C_{G}(\Delta)$ is solvable of bounded derived length. Now $\mathop{\mathrm{Stab}}_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(4)$, and it follows that $\mathop{\mathrm{Stab}}_{G}(\Delta)$ is solvable of bounded derived length. ∎